Solve the following linear system using elimination:
3x+3y=15
5x-5y=15

Answer by checkley77(12844)

3x+3y=15 MULTIPLY BY 5
5x-5y=15 MULTIPLY BY 3 & ADD
15X+15Y=75
15X-15Y=45
------------------
30X=120
X=120/30
X=4 ANSWER.
3*4+3Y=15
12+3Y=15
3Y=15-12
3Y=3
Y=3/3
Y=1 ANSWER.
PROOF:
5*4-5*1=15
20-5=15
15=15